# Large Language Models are Good Spontaneous Multilingual Learners: Is the Multilingual Annotated Data Necessary?

### v1

@@ 7 Conclusion
In this paper, we find that LLMs only trained on translation data without annotated answers are able to get a significant multilingual performance improvement even across a wide range of unseen languages. We conduct experiments on different models, different benchmarks and 20 different languages. Our results indicate that using question translation parallel data can significantly enhance the in-context learning capabilities of LLMs. And these improvements demonstrate excellent model and language generalization. Furthermore, we also conduct comprehensive analysis based on some mechanistic interpretability methods, including logit lens and PCA dimensionality reduction technique. Our work demonstrates the enormous potential of LLMs for efficient multilingual capability improvement. We hope our work can inspire the community to further explore this promising direction for the better multilingual alignment.
@@ 8 Limitations
We aim to draw more attention to the multilingual alignment which is a promising research direction. Despite our work has demonstrated the LLMs’ strong capability of multilingual generalization and the great potential of efficient multilingual enhancement, there are still some limitations waiting for research. Because we investigate the models trained on question translation data without annotated answers in our work, we utilize few-shot learning to help model handle the target task better. Analyzing the models which haven’t been instruction-tuned on the target task properly in a zero-shot setting would further strengthen the conclusions.
Due to the limited resources, we conduct our experiments on different LLM scale from 1.8B to 14B in this work. We are more than willing to verify our conclusions on larger LLMs (70B or larger) if more resources are available in the future. Meanwhile, we mainly utilize automatic translation engine in our work because of the limited resources, while translation data annotated by native speakers would be more accurate.
@@ Appendix B Additional Experimental Implementations
For instruction-tuning process we mentioned above, we use LoRA (rank = 8, $\alpha$ = 16) with 3 epochs, batch_size = 16, learning_rate = 5e-5, val_size = 0.05, lr_scheduler_type = cosine, cutoff_len = 2048 based on the settings of LLaMA-Factory [^3] https://github.com/hiyouga/LLaMA-Factory (Zheng et al., 2024), a widely used and recognized open-source project for LLMs efficient fine-tuning. We use single NVIDIA RTX A6000 48GB or single NVIDIA Tesla V100 SXM2 32GB for training. Training time varies from 4 hours to 10+ hours depending on the language and total instance quantity.
@@
Additionally, we find that not only non-English inputs but also non-English outputs have significant impacts on the model’s performance. For example, for Mistral-7B and emotion classification task, the accuracy on Hindi is 0.5 if we use outputs in Hindi, while the accuracy is 0.816 if the output is "positive" or "negative". This implies that generating content in the target language is another great challenge for LLMs, which is distinct from understanding and solving problems in the corresponding language.
@@ Appendix C Pearson Correlation Coefficient Based on PCA
en-de 0.9711 0.9738
en-fr 0.9798 0.9814
en-hi 0.9140 0.9491
de-fr 0.9823 0.9833
de-hi 0.9489 0.9703
fr-hi 0.9434 0.9664
Table 8: Pearson correlation coefficient of 1 dimension PCA results in layer 20 in Figure 2.
@@
en-de 0.9223 0.9122
en-fr 0.9378 0.9298
en-hi 0.5442 0.7316
de-fr 0.9751 0.9749
de-hi 0.7199 0.8838
fr-hi 0.7109 0.8731
Table 9: Pearson correlation coefficient of 1 dimension PCA results in layer 25 in Figure 2.

### v2

@@ 7 Conclusion
In this paper, we find that LLMs only trained on question translation data without annotated answers are able to get a significant multilingual alignment improvement between English and a wide range of languages, even those unseen during instruction-tuning. We conduct the experiments on different models, different benchmarks and 20 different languages to strengthen our conclusions. Our results indicate that utilizing question alignment significantly enhances the multilingual alignment and the in-context learning capabilities of LLMs. And these improvements demonstrate the excellent model and language generalization. Furthermore, we also conduct comprehensive analysis based on some mechanistic interpretability methods, including logit lens and dimensionality reduction technique. Our work demonstrates the enormous potential of LLMs for efficient multilingual capability improvement. We hope our work can inspire the community to further explore this promising direction for the better multilingual alignment.
@@ 8 Limitations
We aim to draw more attention to the multilingual alignment which is a promising research direction. Despite our work has demonstrated LLMs’ strong capability of multilingual generalization and the great potential of efficient multilingual alignment, there are still some limitations waiting for research. Because we investigate the models trained on the parallel question translation data in our work to eliminate the interference of the task-related data with annotated answers from our analysis of alignment, we utilize few-shot learning to help model handle the target tasks better. Analyzing LLMs’ multilingual alignment in a zero-shot setting properly would further strengthen the conclusions if possible.
Due to the limited resources, we conduct experiments on different LLM scale from 1.8B to 14B in this work. We are willing to verify our conclusions on larger LLMs (70B or larger) if more resources are available in the future. Meanwhile, we mainly utilize automatic translation engine in our work because of the limited resources, while data annotated by native speakers would be more accurate.
@@ Appendix B Additional Experimental Implementations
For instruction-tuning process we mentioned above, we use LoRA (rank = 8, $\alpha$ = 16) with 3 epochs (1 epoch for PAWS to mitigate overfitting), batch_size = 16, learning_rate = 5e-5, val_size = 0.05, lr_scheduler_type = cosine, cutoff_len = 2048 based on the settings of LLaMA-Factory [^5] https://github.com/hiyouga/LLaMA-Factory (Zheng et al., 2024), a widely used and recognized open-source project for LLMs efficient fine-tuning. We use single NVIDIA RTX A6000 48GB or single NVIDIA Tesla V100 SXM2 32GB for training. Training time varies from 4 hours to 10+ hours depending on the language and the total instance quantity.
@@
Additionally, we find that not only non-English inputs but also non-English outputs have significant impacts on the model’s performance. For example, for Mistral-7B and emotion classification task, the accuracy on Hindi is 0.5 if we use outputs in Hindi, while the accuracy is 0.816 if the output is "positive" or "negative". This implies that generating content in the target language is another great challenge for LLMs, which is distinct from understanding and solving problems in the corresponding language. Considering we mainly focus on the language understanding and task solving capabilities, we use English outputs uniformly if it is not specified.
@@ Appendix C Pearson Correlation Coefficient Based on PCA
en-de 0.9727 0.9752
en-fr 0.9804 0.9822
en-hi 0.9268 0.9526
de-fr 0.9825 0.9834
de-hi 0.9564 0.9707
fr-hi 0.9518 0.9674
en-th 0.8727 0.8941
en-sw 0.9501 0.9594
en-ms 0.9552 0.9620
Table 8: Pearson correlation coefficient of 1 dimension PCA results in Mistral-7B layer 20.
@@
en-de 0.9286 0.9275
en-fr 0.9520 0.9447
en-hi 0.5060 0.7942
de-fr 0.9628 0.9754
de-hi 0.6233 0.9102
fr-hi 0.6272 0.9001
en-th 0.6791 0.7484
en-sw 0.2316 0.8514
en-ms 0.7835 0.8448
Table 9: Pearson correlation coefficient of 1 dimension PCA results in Mistral-7B layer 25.
Appendix D Logit Lens and PCA Results for Qwen1.5
We report the logit lens and the PCA results of Qwen1.5-1.8B (total 24 layers) here. In Figure 3, we can find that as we mentioned above, while utilizing logit lens on Qwen1.5, a non-English-centric model, there is no intermediate latent output before generating the output in the target language finally. This indicates that logit lens might not be an effective tool for analyzing the non-English-centric LLMs.
We further report the PCA results in Figure 4, which also indicates a clear similar latent representation pattern for different languages in the non-English-centric LLMs’ intermediate layers. This further reinforces the significance of multilingual alignment, which also provides the basis for the success of question alignment paradigm on Qwen. The Pearson coefficient results reported in Table 10 and Table 11 show the better alignment with English, consisting with the results of Mistral-7B.
Layer 12 Base Trained
en-de 0.9816 0.9820
en-fr 0.9848 0.9850
en-hi 0.9730 0.9742
de-fr 0.9852 0.9855
de-hi 0.9760 0.9769
fr-hi 0.9732 0.9740
en-th 0.9538 0.9547
en-sw 0.9746 0.9763
en-ms 0.9815 0.9823
Table 10: Pearson correlation coefficient of 1 dimension PCA results in Qwen1.5-1.8B layer 12.
Layer 18 Base Trained
en-de 0.7331 0.9176
en-fr 0.6243 0.9303
en-hi 0.3623 0.8589
de-fr 0.7493 0.9477
de-hi 0.5639 0.8965
fr-hi 0.4729 0.8854
en-th 0.2881 0.4566
en-sw -0.0424 0.8592
en-ms 0.4297 0.8648
Table 11: Pearson correlation coefficient of 1 dimension PCA results in Qwen1.5-1.8B layer 18.
(a) English Before
(b) Chinese Before
(c) Japanese Before
(d) English After
(e) Chinese After
(f) Japanese After
Figure 3: Logit lens on Qwen1.5-1.8B in English, Chinese, and Japanese scenarios (languages not in training data). The horizontal axes is the layer num and the vertical axes is the probability. "en" (Green) means the latent English output corresponding to the correct answer in the target language. "en/zh/ja" (Orange) means the correct answer in the target language. "all_possible_out" (Blue) means the probability of all possible outputs in the target language. "all_possible_latent" (Red) means all possible outputs in English.
(a) Layer 6 Before
(b) Layer 12 Before
(c) Layer 18 Before
(d) Layer 24 Before
(e) Layer 6 After
(f) Layer 12 After
(g) Layer 18 After
(h) Layer 24 After
Figure 4: PCA (Principal Component Analysis) on Qwen1.5-1.8B in English, German, French and Hindi scenarios. Before means the base model. After means the trained model. All logits are mapped into the two-dimensional representation. Each point in the plot corresponds to one instance. There is also a similar latent representation pattern for different languages in the intermediate layers while logit lens can’t reveal it.
Appendix E Full Results of SNLI and PAWS
We report the complete results of SNLI and PAWS on 20 different languages in Table 12 and 13 separately. Similar to the emotion classification task, we can see that models instruction-tuned on multilingual translation data significantly outperform the base model, which confirms that our findings have good generalization across different tasks.
Qwen1.5-14B en zh de fr es it nl ja ru sv
base 84.50 83.50 74.17 75.17 81.17 78.67 78.17 51.17 76.83 76.17
zh/de $\Rightarrow$ en 91.83 84.50 83.67 84.50 85.00 83.67 84.50 57.17 81.67 84.33
Qwen1.5-14B sl pl bg no ms is hi th sw bn
base 63.17 67.83 64.33 43.00 75.00 48.17 61.00 69.67 45.83 41.33
Qwen1.5-14B en zh de fr es it nl ja ru sv
base 85.4 78.4 74.0 82.2 78.8 79.4 78.6 75.0 73.0 76.6
zh/es $\Rightarrow$ en 87.8 83.2 81.2 86.0 85.8 83.8 86.0 77.0 82.6 84.4
zh/de $\Rightarrow$ en 87.0 80.8 84.8 84.8 84.6 82.2 86.0 77.0 83.0 84.6
zh/it $\Rightarrow$ en 87.4 80.8 80.4 85.0 83.8 85.2 85.2 77.0 82.2 85.2
sw/hi $\Rightarrow$ en 87.2 78.4 79.6 81.8 84.6 83.2 84.4 76.6 81.0 85.0
Qwen1.5-14B sl pl bg no ms is hi th sw bn
base 70.2 79.8 75.2 78.6 82.0 67.0 73.2 79.4 75.4 65.0
zh/es $\Rightarrow$ en 82.6 80.8 82.0 86.4 86.8 77.4 80.6 82.0 79.2 76.2
zh/de $\Rightarrow$ en 80.8 80.2 82.0 85.8 85.8 77.0 81.4 81.4 80.8 74.2
zh/it $\Rightarrow$ en 81.2 81.6 81.8 86.8 85.6 77.2 79.4 81.2 81.2 74.8
sw/hi $\Rightarrow$ en 82.4 81.8 82.6 87.0 85.8 78.2 80.8 80.4 78.2 77.0
Table 13: Accuracy of Qwen1.5-14B base model and trained models on the PAWS. We report all of the results on 20 languages. The accuracy of randomly choosing is 50.0%. We highlight the best results for every language.